OpenStudy (anonymous):

Can someone help me with Exponential and logarithmic equations,give me examples on how to solve them please

OpenStudy (solomonzelman):

you want some tutoring in general, or do you need help with a particular problem ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{\rm a}=\left(\color{blue}{\rm b}\right)^{\LARGE\color{green}{\rm c}}~~~~~~~_{\Huge\Longrightarrow}^{\rm ~~gives} ~~~~~~\log_{\color{blue}{\LARGE\rm b}}\left(\color{red}{\rm a }\right)=\color{green}{\rm c}}\)

OpenStudy (solomonzelman):

you can prove that by taking log base b on both sides of the a=b^c

OpenStudy (solomonzelman):

bye

OpenStudy (anonymous):

Thank you SolomonZelman